# Statistical Test for Data Analysis Pipeline by Selective Inference

### v1

@@ Abstract
A data analysis pipeline is a structured sequence of processing steps that transforms raw data into meaningful insights by effectively integrating various analysis algorithms. In this paper, we propose a novel statistical test designed to assess the statistical significance of data analysis pipelines. Our approach allows for the systematic development of valid statistical tests applicable to any data analysis pipeline configuration composed of a set of data analysis components. We have developed this framework by adapting selective inference, which has gained recent attention as a new statistical inference technique for data-driven hypotheses. The proposed statistical test is theoretically designed to control the type I error at the desired significance level in finite samples. As examples, we consider a class of pipelines composed of three missing value imputation algorithms, three outlier detection algorithms, and three feature selection algorithms. We confirm the validity of our statistical test through experiments with both synthetic and real data for this class of data analysis pipelines. Additionally, we present an implementation framework that facilitates testing across any configuration of data analysis pipelines in this class without extra implementation costs.
@@ 1 Introduction
In practical data analysis tasks, integrating various types of algorithms is essential to address diverse challenges. For example, in genetic research that seeks to identify genes linked to a specific disease, we often start with preprocessing tasks such as filling in missing values and detecting outliers. Subsequently, we screen for potentially related genes using simple descriptive statistics, followed by more complex machine learning-based feature selection algorithms. Such a systematic sequence of steps designed to analyze data to derive useful insights is referred to as a data analysis pipeline. A data analysis pipeline plays a crucial role in ensuring reproducibility and reliability in data-driven decision-making. Describing the entire data analysis process as a pipeline facilitates precise replication of the analysis, enhancing the consistency of the results. This is especially effective in data-driven scientific research, a well-designed pipeline fortifies the scientific process, allowing researchers to build on each other’s work confidently, advancing knowledge with verified and consistent findings. In this study, as an example, we examine a simple class of data analysis pipelines that integrates various missing-value imputations (MVI) algorithms outlier detection (OD) algorithms, and feature selection (FS) algorithms. Figure 1 displays examples of two such pipelines. The pipeline on the left starts with a mean value imputation algorithm, followed by $L_{1}$ regression based OD algorithm, proceeds with marginal screening to refine feature candidates, and concludes by using two FS algorithms—stepwise feature selection and Lasso—selecting their union as the final features. The pipeline on the right initiates with regression imputation, continues with marginal screening to narrow down feature candidates, uses Cook’s distance for OD, and applies both stepwise FS and Lasso, ultimately choosing the intersection of their results as the final features.
When a data-driven approach is used for critical issues in scientific studies, it is necessary to quantify the reliability of the final results obtained through the pipeline. The goal of this study is to construct a framework for quantifying the statistical reliability of data analysis pipelines within statistical testing framework and quantify the statistical significance in the form of $p$-values. $P$-values provide a measurable criterion to determine whether observed effects are likely due to chance or are statistically significant, making objective decisions regarding the validity of a hypothesis. Furthermore, $p$-values are one of the most commonly used metrics in scientific research to quantify the significance. However, conducting statistical tests on the entire data analysis pipeline is challenging due to the complex interrelations among multiple algorithms. Given the numerous possible combinations of algorithms and their integration, designing individual statistical tests for each pipeline is not practical. Instead, a universal approach to calculating $p$-values for any pipeline configuration is required. To tackle these challenges, we introduce the concept of selective inference (SI) (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014), a novel statistical inference approach that has garnered considerable attention in the past decade. The key concept of SI is to characterize the selection process of hypotheses from the data and calculate the corresponding $p$-value using the sampling distribution conditional on this hypothesis selection process. In this study, we propose an approach based on SI that enables to provide valid $p$-values for any pipeline configuration within a specific class. Furthermore, we develop an implementation framework that facilitates SI for any pipeline configuration within the above class of pipelines without additional implementation cost. The code is available as supplementary material.
@@
Related Works.
@@
Most research on data analysis pipelines is concentrated in the field of software engineering, rather than machine learning (Sugimura and Hartl, 2018; Hapke and Nelson, 2020; Drori et al., 2021). These studies focus on the design, implementation, testing, and maintenance of data analysis pipeline systems to ensure their efficiency, scalability and robustness. Research in AutoML has prompted the creation of frameworks capable of autonomously selecting appropriate algorithms and hyperparameters, effectively automating the construction of data analysis pipelines (Yang et al., 2020; Yakovlev et al., 2020; Karmaker et al., 2021). However, to the best of our knowledge, there is no existing research that systematically discusses the reliability of data analysis pipelines and its automatic generation from a statistical perspective. Resampling techniques such as cross-validation (CV) are commonly used for evaluating the entire data analysis process. However, practical data analysis often includes unsupervised learning tasks such as MVIs and ODs, where resampling cannot be directly applied. Furthermore, dividing the data reduces the sample size, leading to decreased accuracy in hypothesis selection and statistical power.
@@
SI has gained attention as a method of statistical inference for feature selection in linear models (Taylor and Tibshirani, 2015; Fithian et al., 2015). SI for various feature selection algorithms such as marginal screening (Lee and Taylor, 2014), stepwise FS (Tibshirani et al., 2016), and Lasso (Lee et al., 2016) have been explored, and then extended to more complex FS methods, e.g., (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). SI proves valuable not only for FS in linear models but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks such as OD (Chen and Bien, 2020; Tsukurimichi et al., 2021), segmentation (Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024), clustering (Lee et al., 2015; Gao et al., 2022), change-point detection (Duy et al., 2020; Jewell et al., 2022). Moreover, SI is being applied not only to linear models but also to more complex models such as kernel models (Yamada et al., 2018), tree structure models (Neufeld et al., 2022), and CNNs (Duy et al., 2022; Miwa et al., 2023) and Transformers (Shiraishi et al., 2024). The basic form of SI applies to hypothesis selection events characterized as a set of linear inequalities within the data space, with several extensions have been made to apply it to a broader class of selection events (Liu et al., 2018; Tian and Taylor, 2018; Duy and Takeuchi, 2022; Panigrahi et al., 2024). To our knowledge, existing SI research has predominantly concentrated on individual algorithms, with limited exploration into evaluating the statistical significance of outcomes arising from the combination of multiple algorithms.
@@
The summary of our contributions is as follows. First, we develop a statistical test with finite sample validity for data analysis pipelines consisting of a prespecified class of components. Second, we develop a computational and implementation framework that allows for constructing statistical tests across any configuration of pipelines without extra implementation costs. Third, instead of naively applying existing SI, we construct a more powerful version by formulating it as a line search problem and verify its performance through experiments with synthetic and real datasets. Our implementation is available at https://github.com/shirara1016/statistical_test_for_data_analysis_pipeline.
@@ 2 Preliminaries
$\bm{Y}=\bm{\mu}(X)+\bm{\varepsilon},\bm{\varepsilon}\sim N(\bm{0},\Sigma)$ (1)
where $\bm{\mu}(X)\in\mathbb{R}^{n^{\prime}}$ is the unknown true value function, while $\bm{\varepsilon}\in\mathbb{R}^{n^{\prime}}$ is normally distributed noise with covariance matrix $\Sigma$ which is known or estimable from an independent dataset111 We discuss the robustness of the proposed method when the covariance matrix is unknown and the noise deviates from the Gaussian distribution in Appendix E.. Although we do not pose any functional form on the true value function $\bm{\mu}(X)$ for theoretical justification, we consider a case where the true values $\bm{\mu}(X)$ are reasonably approximated by a linear model as long as they are non-outliers. This is a common setting in the field of SI, referred to as the saturated model setting. Furthermore, we denote the response vector with imputed missing values as $\bm{y}^{(+)}\in\mathbb{R}^{n}$. Using the above notations, a data analysis pipeline comprising of MVI, OD, and FS algorithm components is represented as a function
${\mathcal{P}}:\mathbb{R}^{n\times d}\times\mathbb{R}^{n^{\prime}}\ni(\bm{X},\bm{y})\mapsto(\bm{y}^{(+)},{\mathcal{O}},{\mathcal{M}})\in\mathbb{R}^{n}\times 2^{[n]}\times 2^{[d]},$ (2)
@@
$\bm{\beta}^{*}=\left(X_{-{\mathcal{O}},{\mathcal{M}}}^{\top}X_{-{\mathcal{O}},{\mathcal{M}}}\right)^{-}X_{-{\mathcal{O}},{\mathcal{M}}}^{\top}\bm{\mu}^{(+)}_{-{\mathcal{O}}}(X_{-{\mathcal{O}},{\mathcal{M}}}),$ (4)
where $\bm{\mu}^{(+)}_{-{\mathcal{O}}}(X_{-{\mathcal{O}},{\mathcal{M}}})\in\mathbb{R}^{n-|{\mathcal{O}}|}$ indicates a vector of dimension $n-|{\mathcal{O}}|$ that are obtained by providing $X_{-{\mathcal{O}},{\mathcal{M}}}$ to the unknown true function $\bm{\mu}$ with the missing values imputed with the same MVI algorithm. To quantify the statistical significance of the selected features, we consider the following null hypothesis ${\rm H}_{0}$ and the alternative hypothesis ${\rm H}_{1}$:
${\rm H}_{0}:\beta^{*}_{j}=0\text{ v.s. }{\rm H}_{1}:\beta^{*}_{j}\neq 0,j\in{\mathcal{M}},$ (5)
where, with a slight abuse of notation, $\beta^{*}_{j}$ and $\hat{\beta}_{j}$ respectively indicates the element of $\bm{\beta}^{*}$ and ^$\bm{\hat{}}{\beta}$ corresponding to the selected feature $j\in{\mathcal{M}}$.
@@
$f_{\rm MVI}:\{X,\bm{y},{\mathcal{O}},{\mathcal{M}}\}\mapsto\{X,\bm{y}^{+},{\mathcal{O}},{\mathcal{M}}\},$ (6)
@@
$f_{\rm OD}:\{X,\bm{y}^{(+)},{\mathcal{O}},{\mathcal{M}}\}\mapsto\{X,\bm{y}^{(+)},{\mathcal{O}}^{\prime},{\mathcal{M}}\},$ (7)
@@
$f_{\rm FS}:\{X,\bm{y}^{(+)},{\mathcal{O}},{\mathcal{M}}\}\mapsto\{X,\bm{y}^{(+)},{\mathcal{O}},{\mathcal{M}}^{\prime}\},$ (8)
@@
In this study, we consider two cases for pipeline configuration: an option specified by the user and an option determined based on the data. In the first option, the user can select some of the aforementioned data analysis components and specify their own configuration. On the other hand, the second option allows for the selection of the optimal configuration from among multiple pre-defined pipeline configurations based on cross-validation. An important point in the second option is that our statistical test is designed by properly considering the fact that the optimal pipeline configuration has been selected based on the data. For more details on the second option, see §6 and Appendix F.
@@
For the statistical test in (5), it is reasonable to use $\hat{\beta}_{j},j\in{\mathcal{M}}$ as the test statistic. An important point when addressing this statistical test within the SI approach is that the test statistic is represented as a linear function of the observed response vector as $\hat{\beta_{j}}=\bm{\eta}_{j}^{\top}\bm{y},j\in{\mathcal{M}}$, where $\bm{\eta}_{j}\in\mathbb{R}^{n^{\prime}},j\in{\mathcal{M}}$ is a vector that depends on $\bm{y}$ only through the detected outlier set ${\mathcal{O}}$ and the selected feature set ${\mathcal{M}}$ 222 Note that the MVI algorithms considered in this paper depend only on $X$, not on $\bm{y}$. . In SI, this property is utilized to perform statistical inference based on the sampling distribution of the test statistic conditional on ${\mathcal{O}}$ and ${\mathcal{M}}$. More specifically, since $\bm{y}$ follows a normal distribution, it can be derived that the sampling distribution of the test statistic $\hat{\beta}_{j}=\bm{\eta}_{j}^{\top}\bm{y},j\in{\mathcal{M}}$ conditional on ${\mathcal{O}}$, ${\mathcal{M}}$, and the sufficient statistic of the nuisance parameters follows a truncated normal distribution. By computing $p$-values based on this conditional sampling distribution represented as a truncated normal distribution, it is ensured that the type I error can be controlled even in finite samples. For more details on SI, please refer to the explanations in the following sections or literatures such as (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014).
@@ 3 Selective Inference for Data Analysis Pipeline
$\mathbb{P}_{\mathrm{H}_{0}}(p\leq\alpha)=\sum_{\mathcal{M}\in 2^{[p]}}\sum_{\mathcal{O}\in 2^{[n]}}\mathbb{P}_{\mathrm{H}_{0}}(\mathcal{M},\mathcal{O})\mathbb{P}_{\mathrm{H}_{0}}\left(p\leq\alpha\mid\mathcal{M}_{\bm{Y}}=\mathcal{M},\mathcal{O}_{\bm{Y}}=\mathcal{O}\right)=\alpha.$ (12)
@@
To perform statistical hypothesis test based on the conditional sampling distribution in (13), we introduce an additional condition on the sufficient statistic of the nuisance parameter $\mathcal{Q}_{\bm{Y}}$, defined as
@@
This additional conditioning on $\mathcal{Q}(\bm{Y})$ is a standard practice in the SI literature required for computational tractability333 The nuisance component $\mathcal{Q}_{\bm{Y}}$ corresponds to the component $\bm{z}$ in the seminal paper (Lee et al., 2016) (see Sec. 5, Eq. (5.2), and Theorem 5.2) and is used in almost all the SI-related works that we cited. . Based on the additional conditioning on $\mathcal{Q}(\bm{Y})$, the following theorem tells that the conditional $p$-value that satisfies (11) can be derived by using a truncated normal distribution.
@@
is a truncated normal distribution $\mathrm{TN}(\bm{\eta}^{\top}\bm{\mu},\sigma^{2}\|\bm{\eta}\|^{2},\mathcal{Z})$ with the mean $\bm{\eta}^{\top}\bm{\mu}$, the variance $\sigma^{2}\|\bm{\eta}\|^{2}$, and the truncation intervals $\mathcal{Z}$. The truncation intervals $\mathcal{Z}$ is defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{y}},\mathcal{O}_{\bm{a}+\bm{b}z}=\mathcal{O}_{\bm{y}}\right\},\ \bm{a}=\mathcal{Q}_{\bm{y}},\ \bm{b}=\bm{\eta}/\|\bm{\eta}\|^{2}.$ (16)
@@
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{Y})|\geq|T(\bm{y})|\mid\mathcal{M}_{\bm{Y}}=\mathcal{M}_{\bm{y}},\mathcal{O}_{\bm{Y}}=\mathcal{O}_{\bm{y}},\mathcal{Q}_{\bm{Y}}=\mathcal{Q}_{\bm{y}}\right).$ (17)
@@
$\mathbb{P}_{\mathrm{H}_{0}}\left(p_{\mathrm{selective}}\leq\alpha\mid\mathcal{M}_{\bm{Y}}=\mathcal{M}_{\bm{y}},\mathcal{O}_{\bm{Y}}=\mathcal{O}_{\bm{y}}\right)=\alpha,\ \forall\alpha\in(0,1).$ (18)

### v2

# Statistical Test for Feature Selection Pipelines by Selective Inference

@@ Abstract
A data analysis pipeline is a structured sequence of steps that transforms raw data into meaningful insights by integrating various analysis algorithms. In this paper, we propose a novel statistical test to assess the significance of data analysis pipelines in feature selection problems. Our approach enables the systematic development of valid statistical tests applicable to any feature selection pipeline composed of predefined components. We develop this framework based on selective inference, a statistical technique that has recently gained attention for data-driven hypotheses. As a proof of concept, we consider feature selection pipelines for linear models, composed of three missing value imputation algorithms, three outlier detection algorithms, and three feature selection algorithms. We theoretically prove that our statistical test can control the probability of false positive feature selection at any desired level, and demonstrate its validity and effectiveness through experiments on synthetic and real data. Additionally, we present an implementation framework that facilitates testing across any configuration of these feature selection pipelines without extra implementation costs.
@@ 1 Introduction
In practical data-driven decision-making tasks, integrating various types of data analysis steps is crucial for addressing diverse challenges. For instance, in genetic research aimed at identifying genes linked to a specific disease, the process often begins with preprocessing tasks such as filling in missing values and detecting outliers. This is followed by screening for potentially related genes using simple descriptive statistics and then applying more complex machine learning-based feature selection algorithms. Such a systematic sequence of steps designed to analyze data and derive useful insights is known as a data analysis pipeline, which plays a key role in ensuring the reproducibility and reliability of data-driven decision-making.
In this study, as an example of data analysis pipelines, we consider a class of feature selection pipelines that integrates various missing-value imputations (MVI) algorithms, outlier detection (OD) algorithms, and feature selection (FS) algorithms. Figure 1 shows examples of two such pipelines. The pipeline on the left starts with a mean value imputation algorithm, followed by $L_{1}$ regression based OD algorithm, proceeds with marginal screening to refine feature candidates, and concludes by using two FS algorithms—stepwise feature selection and Lasso—selecting their union as the final features. The pipeline on the right initiates with regression imputation, continues with marginal screening to narrow down feature candidates, uses Cook’s distance for OD, and applies both stepwise FS and Lasso, ultimately choosing the intersection of their results as the final features.
@@
When a data-driven approach is used for high-stakes decision-making tasks such as medical diagnosis, it is crucial to quantify the reliability of the final results by considering all steps in the pipeline. The goal of this study is to develop a statistical test for a specific class of feature selection pipelines, allowing the statistical significance of features obtained through the pipeline to be properly quantified in the form of $p$-values. The first technical challenge in achieving this is the need to appropriately account for the complex interrelations between pipeline components to determine the overall statistical significance. The second challenge is to develop a universal framework capable of performing statistical tests on arbitrary pipelines (within a given class) rather than creating individual tests for each pipeline.
To address these challenges, we introduce the concept of selective inference (SI) (Taylor and Tibshirani, 2015; Fithian et al., 2015; Lee and Taylor, 2014), a novel statistical inference approach that has gained significant attention over the past decade. The core idea of SI is to characterize the process of selecting hypotheses from the data and calculate the corresponding $p$-values using the sampling distribution, conditional on this selection process. We propose an approach based on SI that provides valid $p$-values for any feature selection pipeline configuration within the aforementioned class. We also introduce a modular implementation framework that supports SI for any pipeline configuration within this class without requiring additional implementation efforts [^1] See our 2-minute video in the supplementary material demonstrating this implementation framework.. Specifically, with our framework, the statistical significance of features from any pipeline in this class can be quantified as valid $p$-values when used in a linear model, with no extra implementation required beyond specifying the pipeline.
We note that our long-term goal beyond this current study is to ensure the reproducibility of data-driven decision-making by accounting for the entire pipeline from raw data to the final results, with the current study on a class of feature selection pipelines serving as a proof of concept for that goal.
Related Work.
@@
Most research on data analysis pipelines is concentrated in the field of software engineering rather than machine learning (Sugimura and Hartl, 2018; Hapke and Nelson, 2020; Drori et al., 2021), with a primary focus on the design, implementation, testing, and maintenance of pipeline systems to ensure efficiency, scalability, and robustness. Meanwhile, AutoML has emerged as a related area where researchers are automating the construction of these pipelines, and many companies have developed tools for this purpose (Microsoft, 2018; Amazon, 2019; Google, 2021). However, to the best of our knowledge, there is no existing studies that systematically discusses the reliability of data analysis pipelines. Resampling techniques, such as cross-validation (CV), are commonly used to evaluate the entire data analysis process. However, practical data analysis often includes unsupervised learning tasks like MVIs and ODs, where resampling cannot be used to accurately evaluate the reliability of the entire pipeline. Additionally, dividing the data reduces the sample size, leading to decreased accuracy in hypothesis selection and statistical power.
@@
SI has gained attention as a statistical inference method for feature selection in linear models (Taylor and Tibshirani, 2015; Fithian et al., 2015). It has been applied to various feature selection algorithms such as marginal screening (Lee and Taylor, 2014), stepwise FS (Tibshirani et al., 2016), and Lasso (Lee et al., 2016), and extended to more complex methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). SI is valuable not only for FS in linear models but also for inference across various data-driven hypotheses, including tasks like OD (Chen and Bien, 2020; Tsukurimichi et al., 2021), segmentation (Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024), clustering (Lee et al., 2015; Gao et al., 2022), and change-point detection (Duy et al., 2020; Jewell et al., 2022). The core idea of selective inference (SI) is to perform statistical inference using a distribution conditioned on events of hypothesis selection, with the technical challenge being the characterization of various event selections for different tasks. While studies on statistical inference (SI) for various tasks are being conducted, existing research is limited to single tasks, and how to perform inference when integrating multiple tasks into a pipeline remains an open question. Furthermore, existing implementations of SI are developed individually for each task, and there is no unified framework for implementing SI.
@@
Our contributions in this study are threefold. First, we develop a statistical test for feature selection pipelines composed of various configurations of missing value imputation (MVI), outlier detection (OD), and feature selection (FS) components, based on the SI framework. Second, this study represents the first application of SI to inference on a combination of multiple analysis components in a unified, systematic manner. Finally, we establish a computational and implementation framework, available as supplementary material, that facilitates the construction of statistical tests across any pipeline configuration without additional implementation costs. Our supplementary material is available at https://github.com/shirara1016/statistical_test_for_feature_selection_pipelines.
@@ 2 Preliminaries
$\bm{Y}=\bm{\mu}(X)+\bm{\varepsilon},\ \bm{\varepsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}I_{n^{\prime}}),$ (1)
where $\bm{\mu}(X)\in\mathbb{R}^{n^{\prime}}$ is the unknown true value function, while $\bm{\varepsilon}\in\mathbb{R}^{n^{\prime}}$ is independently normally distributed with variance $\sigma^{2}$ which is known or estimable from an independent dataset222 We discuss the robustness of the proposed method when the variance is unknown and the noise deviates from the Gaussian distribution in Appendix E.. Although we do not pose any functional form on the true value function $\bm{\mu}(X)$ for theoretical justification, we consider a case where the true values $\bm{\mu}(X)$ are reasonably approximated by a linear model as long as they are non-outliers. This is a common setting in the field of SI, referred to as the saturated model setting. Furthermore, we denote the response vector with imputed missing values as $\bm{y}^{(+)}\in\mathbb{R}^{n}$. Using the above notations, a data analysis pipeline comprising of MVI, OD, and FS algorithm components is represented as a function
$\mathcal{P}:\mathbb{R}^{n\times d}\times\mathbb{R}^{n^{\prime}}\ni(X,\bm{y})\mapsto(\bm{y}^{(+)},\mathcal{O},\mathcal{M})\in\mathbb{R}^{n}\times 2^{[n]}\times 2^{[d]},$ (2)
@@
$\bm{\beta}^{\ast}=\left(X_{-\mathcal{O},\mathcal{M}}^{\top}X_{-\mathcal{O},\mathcal{M}}\right)^{-}X_{-\mathcal{O},\mathcal{M}}^{\top}\bm{\mu}^{(+)}_{-\mathcal{O}}(X_{-\mathcal{O},\mathcal{M}}),$ (4)
where $\bm{\mu}^{(+)}_{-\mathcal{O}}(X_{-\mathcal{O},\mathcal{M}})\in\mathbb{R}^{n-|\mathcal{O}|}$ indicates a vector of dimension $n-|\mathcal{O}|$ that are obtained by providing $X_{-\mathcal{O},\mathcal{M}}$ to the unknown true function $\bm{\mu}$ with the missing values imputed with the same MVI algorithm. To quantify the statistical significance of the selected features, we consider the following null hypothesis $\mathrm{H}_{0}$ and the alternative hypothesis $\mathrm{H}_{1}$:
$\mathrm{H}_{0}:\beta^{\ast}_{j}=0\ \text{v.s.}\ \mathrm{H}_{1}:\beta^{\ast}_{j}\neq 0,\ j\in\mathcal{M},$ (5)
where, with a slight abuse of notation, $\beta^{\ast}_{j}$ and $\hat{\beta}_{j}$ respectively indicates the element of $\bm{\beta}^{\ast}$ and $\hat{\bm{\beta}}$ corresponding to the selected feature $j\in\mathcal{M}$.
@@
$f_{\mathrm{MVI}}:\{X,\bm{y},\mathcal{O},\mathcal{M}\}\mapsto\{X,\bm{y}^{+},\mathcal{O},\mathcal{M}\},$ (6)
@@
In this study, we consider two cases for pipeline configuration: an option specified by the user and an option determined based on the data. In the first option, the user can select some of the aforementioned data analysis components and specify their own configuration. On the other hand, the second option allows for the selection of the optimal configuration from among multiple pre-defined pipeline configurations based on CV. An important point in the second option is that our statistical test is designed by properly considering the fact that the optimal pipeline configuration has been selected based on the data [^3] As stated in §1, CV cannot be used for an accurate evaluation of a pipeline when it includes unsupervised learning components such as MVI or OD. However, it is possible to compute a valid $p$-value for a pipeline selected by CV if we properly consider the CV-based pipeline selection as part of the selection event for SI.. For more details on the second option, see §6 and Appendix F.
@@
For the statistical test in (5), it is reasonable to use $\hat{\beta}_{j},j\in\mathcal{M}$ as the test statistic. An important point when addressing this statistical test within the SI approach is that the test statistic is represented as a linear function of the observed response vector as $\hat{\beta}_{j}=\bm{\eta}_{j}^{\top}\bm{y},j\in\mathcal{M}$, where $\bm{\eta}_{j}\in\mathbb{R}^{n^{\prime}},j\in\mathcal{M}$ is a vector that depends on $\bm{y}$ only through the detected outlier set $\mathcal{O}$ and the selected feature set $\mathcal{M}$ 444 Note that the MVI algorithms considered in this paper depend only on $X$, not on $\bm{y}$. . In SI, this property is utilized to perform statistical inference based on the sampling distribution of the test statistic conditional on $\mathcal{O}$ and $\mathcal{M}$. More specifically, since $\bm{y}$ follows a normal distribution, it can be derived that the sampling distribution of the test statistic $\hat{\beta}_{j}=\bm{\eta}_{j}^{\top}\bm{y},j\in\mathcal{M}$ conditional on $\mathcal{O}$, $\mathcal{M}$, and the sufficient statistic of the nuisance parameters follows a truncated normal distribution. By computing $p$-values based on this conditional sampling distribution represented as a truncated normal distribution, it is ensured that the type I error can be controlled even in finite samples. For more details on SI, please refer to the explanations in the following sections or literatures such as Taylor and Tibshirani (2015); Fithian et al. (2015); Lee and Taylor (2014).
@@ 3 Selective Inference for Data Analysis Pipeline
$\mathbb{P}_{\mathrm{H}_{0}}(p\leq\alpha)$ (12)
[…]
$=\alpha.$ (16)
@@
To conduct statistical hypothesis testing based on the conditional sampling distribution in (17), we introduce an additional condition on the sufficient statistic of the nuisance parameter $\mathcal{Q}_{\bm{Y}}$, defined as
@@
This additional conditioning on $\mathcal{Q}(\bm{Y})$ is a standard practice in the SI literature required for computational tractability555 The nuisance component $\mathcal{Q}_{\bm{Y}}$ corresponds to the component $\bm{z}$ in the seminal paper (Lee et al., 2016) (see Sec. 5, Eq. (5.2), and Theorem 5.2) and is used in almost all the SI-related works that we cited. . Based on the additional conditioning on $\mathcal{Q}(\bm{Y})$, the following theorem tells that the conditional $p$-value that satisfies (11) can be derived by using a truncated normal distribution.
@@
is truncated normal distribution $\mathrm{TN}(\bm{\eta}^{\top}\bm{\mu},\sigma^{2}\|\bm{\eta}\|^{2},\mathcal{Z})$ with the mean $\bm{\eta}^{\top}\bm{\mu}$, the variance $\sigma^{2}\|\bm{\eta}\|^{2}$, and the truncation intervals $\mathcal{Z}$. The truncation intervals $\mathcal{Z}$ is defined as
$\mathcal{Z}=\left\{z\in\mathbb{R}\mid\mathcal{M}_{\bm{a}+\bm{b}z}=\mathcal{M}_{\bm{y}},\mathcal{O}_{\bm{a}+\bm{b}z}=\mathcal{O}_{\bm{y}}\right\},$ (20)
$\bm{a}=\mathcal{Q}_{\bm{y}},\ \bm{b}=\bm{\eta}/\|\bm{\eta}\|^{2}.$ (21)
@@
$p_{\mathrm{selective}}=\mathbb{P}_{\mathrm{H}_{0}}\left(|T(\bm{Y})|\geq|T(\bm{y})|\middle|\begin{gathered}\mathcal{M}_{\bm{Y}}=\mathcal{M}_{\bm{y}},\\ \mathcal{O}_{\bm{Y}}=\mathcal{O}_{\bm{y}},\\ \mathcal{Q}_{\bm{Y}}=\mathcal{Q}_{\bm{y}}\end{gathered}\right).$ (22)
@@
$\mathbb{P}_{\mathrm{H}_{0}}\left(p_{\mathrm{selective}}\leq\alpha\middle|\begin{gathered}\mathcal{M}_{\bm{Y}}=\mathcal{M}_{\bm{y}},\\ \mathcal{O}_{\bm{Y}}=\mathcal{O}_{\bm{y}}\end{gathered}\right)=\alpha,\ \forall\alpha\in(0,1).$ (23)